# Limits to scalable evaluation at the frontier: LLM as Judge won’t beat twice the data

### v1

@@ 3.6 Balanced agreement rate as an alternative
For any value of $\operatorname*{\mathrm{BA}}(m)$, we have:
Theorem 9 is proven in Appendix C.3. It also provides another upper bound on sample efficiency gains from access to $\tilde{s}$: Whenever $\operatorname*{\mathrm{BA}}(m)$ is close to half, access to $\tilde{s}$ can not improve sample efficiency by much and the sample efficiency factor $\tau$ is close to one. As a corollary, we show that whenever the proxy has reasonable balanced accuracy, the minimum of the true positive rate $p$ and true negative rate $q$ upper bounds the squared correlation $\rho^{2}$:
Corollary 10.
Whenever $\operatorname*{\mathrm{BA}}(m)\geq 0.5$, we have
$\rho(s(m),\tilde{s}(m))^{2}\leq\min\{p(m),q(m)\}.$
Proof.
4 Experiments
Our experiments focus on two settings: Multiple-choice question answering and chatbot evaluation. For the first, we evaluate the accuracy of models on MMLU (Hendrycks et al., 2021), a multiple-choice question answering benchmark that aims to measure the world knowledge of language models. We aggregate all subtask test sets of MMLU to obtain a single test set $T$, and measure the performance of a model $m$ on MMLU in terms of accuracy $\frac{1}{|T|}\sum_{x}\mathbb{I}(m(x)=y(x))$. Here $\mathbb{I}$ is the indicator function, $m(x)$ is the model’s answer and $y(x)$ is the correct answer. For model-based evaluations using model $\tilde{m}$ as a judge, we simply replace the correct answer $y(x)$ by the judge’s prediction $\tilde{m}(x)$ and calculate accuracy based on the modified answers.
Figure 2: Best possible sample efficiency factor $\tau_{\max}$ using different judges (colors) evaluating different models (x-ticks) on MMLU. Error bars show $90\%$ confidence intervals.
In the second setting, we evaluate model performance on MT-Bench (Zheng et al., 2024), an arena-style benchmark designed to evaluate the ability of models to follow instructions. The benchmark comes with results for six models. For each pair of these models $m$ and $m^{\prime}$, both models are queried with the same prompt $x$. The evaluator then decides, which model’s response was better. For each model $m$, we calculate its win rate averaged over all triples of the form $(m,m^{\prime},x)$, for other models in the benchmark $m^{\prime}$ and prompts $x$. For more details, consider Appendix B.
Figure 3: Best possible sample efficiency factor $\tau_{\max}$ using GPT-4-as-a-Judge, evaluating different models on MT-Bench. Error bars show $90\%$ confidence intervals.
Figures 2 and 3 show the optimal sample efficiency factor $\tau_{\max}$ on MMLU and MT-Bench respectively. The sample efficiency factor $\tau_{\max}$ consistently stays below the value of two implied by Corollary 7, even when slightly stronger models like GPT-4 are used to evaluate weaker models like LLama3-70B. We only get a larger sample efficiency factor $\tau_{\max}$ when current flagship models are used to judge the significantly worse LLama2-7b on MMLU. Notably, $\tau_{\max}$ is often substantially smaller than two, especially on MMLU. This highlights that Corollary 7 is a best-case upper bound rather than a guarantee that sample efficiency factors close to two are actually achievable in practice.
Going beyond binary evaluations.
We also experiment with incorporating model uncertainty into the proxy score on MMLU. We define the proxy score as the (renormalized) probability the judge model $\tilde{m}$ assigned to model $m$’s answer $\tilde{s}(m)=\operatorname*{\mathbb{P}}_{\tilde{m}(x)}\{m(x)\}$. While this non-binary score $\tilde{s}$ breaks the assumptions of Theorem 5, we expect PPI to remain practically optimal. Figure 4 shows that for LLama3.1-405B as judge, this approach increases the sample efficiency of PPI. However, when evaluating the strongest models, the sample efficiency factor $\tau(\hat{\theta}_{\lambda^{*}}^{PP})$ still remains below two.
@@ 4 Experiments
Figure 4: Data efficiency factor $\tau(\hat{\theta}_{\lambda^{*}}^{PP})$ in the $N\to\infty$ limit, using LLama3.1-405B-as-a-Judge on MMLU, with accuracy scoring vs soft accuracy scoring. Error bars show $90\%$ confidence intervals.
@@ 6 Discussion
First, our results pertain to uniform sampling rather than more sophisticated sampling strategies. Approaches like stratified PPI (Fisch et al., 2024) might be able to obtain a somewhat better sample efficiency factor, for example by breaking the assumptions of Theorem 6 per-stratum. Second, Theorem 6 only holds if the proxy $\tilde{s}$ is less capable at predicting $s$ than the evaluated model $m$ is at its task. Thus, for tasks in which evaluation is substantially easier than the task itself, or tasks for which data to train a specialized evaluator is abundant, gains in sample efficiency of more than a factor two are possible even when evaluating models at the frontier. In addition, weaker models might find their way into many applications due to cost saving reasons, such that efficient evaluations of these models are still valuable. That said, Theorem 5 still applies per-stratum, as well as for weaker models. Correspondingly, variance reduction is still limited by the squared correlation $\rho(s(m),\tilde{s}(m))^{2}$.
@@ A.1 Sample efficiency factor for strictly better classifiers
A.2 Estimating the gap
@@ Appendix B Additional details on experiments
High agreement is not necessary for ranking.

### v2

@@ 3.7 Experiments
For any proxy score $\tilde{s}$ with $\operatorname*{\mathrm{SO}}(R(\tilde{s}))\leq b$, we have $\rho^{2}(s,\tilde{s})\leq 0.5$. Correspondingly, the sample effficiency of PPI is bounded: $\tau(\hat{\theta}^{PP}_{\lambda^{*}})\leq 2$.
We prove more general versions of Proposition 9 and Theorem 10 in Appendix C.3. In these, we allow for the judge to make $\epsilon$ times as many mistakes as the evaluated model for $\epsilon\leq 1$ rather than being strictly worse. In that case, we obtain a maximal sample efficiency factor $\tau(\hat{\theta}^{PP})$ of $\frac{2}{\epsilon}$.
Experiments on MMLU
Using the prompt format from HELM, we extract LLama3.1-405B’s next-token predictions $\tilde{p}_{x}$ for all questions $x$ in MMLU. From these, we take the probabilities $\tilde{p}_{x}(t)$ corresponding to the four tokens $t\in\{A,B,C,D\}\eqqcolon T$ representing the answers to question $x$. We then renormalize them to $\operatorname*{\mathbb{P}}_{m^{\prime}(x)}(t)=(\tilde{p}_{x}(t))/(\sum_{t^{\prime}\in T}\tilde{p}_{x}(t^{\prime}))$ and define the proxy score $\tilde{s}(x,m)$ as the probability $\operatorname*{\mathbb{P}}_{m^{\prime}(x)}(m(x))$ assigned to the evaluated model’s answer $m(x)$, as described above.
Figure 4 shows the sample efficiency factor $\tau(\hat{\theta}^{PP}_{\lambda^{*}})$ for PPI in the limit of infinite unlabeled data $N\to\infty$. As expected, using the non-binary proxy consistently improves sample efficiency. However, as suggested by Theorem 10, the sample efficiency factor $\tau(\hat{\theta}_{\lambda^{*}}^{PP})$ remains below two when we use LLama3.1-405B to evaluate the stronger Claude 3.5. As in the previous experiments, the same is true for evaluating slightly weaker models such as GPT-4 or Gemini 1.5. This shows that our main claim of ”no sample efficiency gains of more than a factor two at the frontier” is robust. That said, unlike in the binary case, the factor two is already (slightly) exceeded for LLaMa3-70B (which is 5% less accurate than LLama3.1-405B), rather than just for the much weaker LLaMa2-7B.
@@
Figure 4: Sample efficiency factor $\tau(\hat{\theta}_{\lambda^{*}}^{PP})$ for PPI in the $N\to\infty$ limit, using LLama3.1-405B-as-Judge on MMLU, with binary and non-binary scores. Error bars: $90\%$ confidence intervals. Non-binary scores improve the sample efficiency factor $\tau(\hat{\theta}_{\lambda^{*}}^{PP})$, but it stays below two at the frontier.
@@ 6 Discussion
First, our results pertain to uniform sampling rather than more sophisticated sampling strategies. Approaches like stratified PPI (Fisch et al., 2024) might be able to obtain a somewhat better sample efficiency factor, for example by breaking the assumptions of Theorem 6 per-stratum. That said, Theorem 5 still applies per-stratum. Empirical results presented in Appendix B.3 suggest that at the frontier, stratified PPI rarely improves sample efficiency by more than a factor of two, compared to standard stratified sampling. Second, Theorem 6 only holds if the proxy $\tilde{s}$ is less capable at predicting $s$ than the evaluated model $m$ is at its task. Thus, for tasks in which evaluation is substantially easier than the task itself, or tasks for which data to train a specialized evaluator is abundant, gains in sample efficiency of more than a factor two are possible even when evaluating models at the frontier. In addition, weaker models might find their way into many applications due to cost saving reasons, such that efficient evaluations of these models are still valuable.
@@ A.1 Sample efficiency factor for strictly better classifiers
A.2 Balanced agreement rate as an alternative
Intuitively, high agreement fails to constrain the squared correlation $\rho^{2}$ because predictors without any signal, such as constant predictors, can achieve high agreement. This is analogous to how accuracy can be misleading for imbalanced classification problems. In classification, this is solved by focusing on balanced accuracy instead. By analogy, we introduce the balanced agreement of $\tilde{s}$ with $s$ as
$\operatorname*{\mathrm{BA}}(m)\coloneqq\frac{q(m)+p(m)}{2}.$
Theorem 12.
For any value of $\operatorname*{\mathrm{BA}}(m)$, we have:
Theorem 12 is proven in Appendix C.4. It also provides another upper bound on sample efficiency gains from access to $\tilde{s}$: Whenever $\operatorname*{\mathrm{BA}}(m)$ is close to half, access to $\tilde{s}$ can not improve sample efficiency by much and the sample efficiency factor $\tau_{\max}=\frac{1}{1-\rho^{2}}$ is close to one. As a corollary, we show that whenever the proxy has reasonable balanced accuracy, the minimum of the true positive rate $p$ and true negative rate $q$ upper bounds the squared correlation $\rho^{2}$:
Corollary 13.
Whenever $\operatorname*{\mathrm{BA}}(m)\geq 0.5$, we have
$\rho(s(m),\tilde{s}(m))^{2}\leq\min\{p(m),q(m)\}.$
Proof.
Figure 5 shows actual value of the maximal sample efficiency and its lower and upper bounds on $\tau_{\max}$ derived from Theorem 12 for a variety of models. In line with our theoretical results, $\tau_{\max}$ lies consistently between the lower and upper bounds. Interestingly, the lower bound appear to be considerably tighter than the upper bound, that is, $\tau_{\max}$ is much closer to the lower bound than the upper bound throughout the experiments.
Figure 5: Best possible sample efficiency factor $\tau_{\max}$ and lower/upper bounds based on the balanced accuracy $\operatorname*{\mathrm{BA}}$ for different judges (color) and evaluated models (x-ticks).
A.3 Estimating the gap
@@ Appendix B Additional details on experiments
B.1 High agreement is not necessary for ranking.
@@
B.2 Additional results on TruthfulQA
Figure 7 shows additional results on TruthfulQA. As on MMLU, we extract model answers to the Q&A prompts from HELM [^2] We accessed the data on 14.11.2024, but it appears as if the leaderboard has not been updated to include recent models such as GPT-4. (Liang et al., 2023) and define the proxy score $\tilde{s}$ using the judge’s answers $m(x)$ as described in Section 2. We use the top three models in the leaderboard as judges, and consider evaluations of these as well as some worse-performing models. As predicted by our theory, the sample efficiency factor $\tau_{\max}$ is consistently below two when weaker models are used to evaluate stronger ones. Beyond that, the sample efficiency factor $\tau_{\max}$ even stays below two when we use the strongest model we considered (Palmyra-x) to judge the weakest (LLaMa2 7B).
Figure 7: Best possible sample efficiency factor $\tau_{\max}$ using different judges (colors) evaluating different models (x-ticks) on TruthfulQA. Error bars show $90\%$ confidence intervals.
B.3 Additional results on stratification
We consider a debaising method for more sophisticated sampling strategies, namely stratified PPI (Fisch et al., 2024). Note that efficiency improvements from stratification occur independently from the model-as-judge regime. In order to disentangle gains from stratification from the gains of (stratified) PPI (Fisch et al., 2024), we therefore focus on the per-stratum sample efficiency factor $\tau_{\max}$. Theorem 6 still holds per-stratum when applying stratified PPI, as long as the assumptions are valid for every stratum. It might however happen that the assumptions do not hold for every stratum, even if they are true globally: A judge model that is weaker than the evaluated model on average could still be stronger on certain strata, thus achieving a sample efficiency factor larger than two (on these strata).
Figure 8 suggests that this might be rare in practice: It shows the per-stratum sample efficiency factors $\tau_{\max}$ on MMLU, using subtasks as strata. Using GPT-4 to judge the stronger Claude (Figure 8(a)) and LLaMa3.1 405B (Figure 8(b)), the sample efficiency factors $\tau_{\max}$ varies significantly across strata. They are larger for most strata than the total (non-stratified) sample efficiency factor. However they consistently stay below two. Even when using the stronger LLaMa3.1 405B (Figure 8(d)) and Claude (Figure 8(c)) to judge GPT-4, sample efficiency gains above two remain rare: When only observe them in 6 out of 57 (Claude) and 1 out of 57 (LLaMa) cases.
(a) GPT-4 judging Claude
(b) GPT-4 judging LLaMa3.1 405B
(c) Claude judging GPT-4
(d) LLaMa3.1 405B judging GPT-4
Figure 8: Per-stratum sample efficiency factor $\tau_{\max}$ on MMLU. The strata are the 57 MMLU subtasks, in alphabetical order.
@@ C.3 Theorem 10
Instead of Theorem 10, we prove a stronger version that allows for the judge to be better than the evaluated model. Concretely, the soft error rate of the judge $(1-\operatorname*{\mathrm{SO}}(R(\tilde{s})))$ is allowed to be smaller than the soft error rate of the evaluated model $1-b$ by a factor $\epsilon<1$. In that case, our upper bound on $\tau$ becomes $\frac{2}{\epsilon}$.
Theorem 15.
For any proxy score $\tilde{s}$ with $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{SO}}(R(\tilde{s})))$, we have $\rho^{2}(s,\tilde{s})\leq 1-\frac{\epsilon}{2}$. Correspondingly, the sample effficiency of PPI is bounded: $\tau(\hat{\theta}^{PP}_{\lambda^{*}})\leq\frac{2}{\epsilon}$.
Proof.
First, we note that the because the Mean squared error/Brier score is a proper scoring rule (Gneiting & Raftery, 2007), the Bayes-optimal predictor $\operatorname*{\mathrm{R}}(\tilde{s})$ minimizes the mean squared error $\operatorname*{\mathrm{MSE}}(g(\tilde{s}),s)$ for all post-processing functions $g$. At the same time, we have that the squared correlation $\rho^{2}(s,\tilde{s})=R^{2}(s,f(\tilde{s}))=1-\frac{\operatorname*{\mathrm{MSE}}(s,f(\tilde{s}))}{\operatorname*{\mathrm{Var}}(s)}$ falls monotonously in $\operatorname*{\mathrm{MSE}}(s,f(\tilde{s}))$, where $f$ is the affine transformation with the smallest value of $\operatorname*{\mathrm{MSE}}(s,f(\tilde{s}))$ (Larsen & Marx, 2005). This means that the bayes-optimal predictor $\operatorname*{\mathrm{R}}(\tilde{s})$ fulfills $\rho^{2}({s,\operatorname*{\mathrm{R}}(\tilde{s})})\geq\rho^{2}(s,{\tilde{s}})$. As the Bayes-optimal predictor is calibrated, it is sufficient to prove our upper bound for calibrated scores $\tilde{s}$ (i.e. scores $\tilde{s}$ such that $\operatorname*{\mathbb{P}}(s|\tilde{s}=\gamma)=\gamma$ for all $\gamma\in[0,1]$.
In the following, we assume $\tilde{s}$ to be calibrated. This means that
$\operatorname*{\mathrm{MSE}}(s,\tilde{s})=\operatorname*{\mathbb{E}}(s-\tilde{s})^{2}=\int_{0}^{1}((1-\tilde{s})^{2}\tilde{s}+\tilde{s}^{2}(1-\tilde{s}))p(\tilde{s})d\tilde{s}=\int_{0}^{1}(1-\tilde{s})\tilde{s}p(\tilde{s})d\tilde{s},$
where we use that because of calibration, conditional on a prediction $\tilde{s}$, $s$ is one and the squared error is $(1-\tilde{s})^{2}$ with probability $\tilde{s}$. Similarly, the squared error is $\tilde{s}^{2}$ with probability $(1-\tilde{s})$.
At the same time, the soft accuracy of $\tilde{s}$ equals
[…]
for calibrated scores $\tilde{s}$, as $s=1$ yields a score of $\tilde{s}$ and happens with probability $\tilde{s}$, and $s=0$ happens with probability $(1-\tilde{s})$ and gets a score of $1-\tilde{s}$.
By the linearity of the integral, this implies that
$\operatorname*{\mathrm{SO}}=1-2\operatorname*{\mathrm{MSE}}$
or
$\operatorname*{\mathrm{MSE}}=\frac{1-\operatorname*{\mathrm{SO}}}{2}.$
As $\tilde{s}$ is bayes-optimal with respect to itself, the optimal affine transformation in terms of $MSE$ is just the identity. Thus, we have that
$\rho^{2}(s,\tilde{s})=R^{2}(\tilde{s},s)=1-\frac{\operatorname*{\mathrm{MSE}}}{\operatorname*{\mathrm{Var}}(s)}=1-\frac{\frac{1-\operatorname*{\mathrm{SO}}}{2}}{b(1-b)}=1-\frac{1-\operatorname*{\mathrm{SO}}}{2b(1-b)}.$
This is decreasing in $1-\operatorname*{\mathrm{SO}}$ and thus maximized at the minial value of $1-\operatorname*{\mathrm{SO}}$, which equals $\epsilon(1-b)$ by assumption. Inserting yields
$\rho^{2}(\tilde{s},s)\leq 1-\frac{\epsilon(1-b)}{2b(1-b)}=1-\frac{\epsilon}{2b}.$
This is increasing in $b$ and thus maximized at $b=1$, where we have
$\rho^{2}(\tilde{s},s)\leq 1-\frac{\epsilon(1-b)}{2b(1-b)}=1-\frac{\epsilon}{2}.$
Together with a generalized version of Proposition 9, Theorem 15 allows us to generalize theorem 6 from the paper:
Theorem 16.
Assume that $b\geq 0.5$. Then for any $0<\epsilon\leq 1$, and any binary proxy $\tilde{s}$ such that $\operatorname*{\mathrm{AG}}\geq 0.5$ and $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{AG}})$, we have $\rho^{2}\leq 1-\frac{\epsilon}{2}$. Correspondingly, the sample efficiency factor $\tau_{\max}$ is at most $\frac{2}{\epsilon}$.
Proof.
We first state the generalized version of Proposition 9.
Proposition 17.
Assume that $b\geq 0.5$. Then for any $0<\epsilon\leq 1$, take any binary proxy $\tilde{s}$ such that $\operatorname*{\mathrm{AG}}\geq 0.5$ and $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{AG}})$. Then, the recalibrated proxy $R(\tilde{s})$ fulfills $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{SO}})$.
We now simply apply Theorem 15. By Proposition 9, the recalibrated version $R(\tilde{s})$ fulfills $\epsilon(1-b)\leq(1-\operatorname*{\mathrm{SO}})$, yielding the desired bound. In the following, we prove Proposition 17:
Proof.
We first characterize the recalibrated proxy $R(\tilde{s})$ and its relationship with the real score $s$: $R(\tilde{s})$ equals $\operatorname*{\mathbb{P}}(s=1|\tilde{s}=1)$ whenever $\tilde{s}=1$ and $\operatorname*{\mathbb{P}}(s=1|\tilde{s}=0)$ whenever $\tilde{s}=0$. This means that
$R(\tilde{s})=\begin{cases}\frac{bp}{bp+(1-b)(1-q)}\eqcolon x_{1}\ w.p.\ bp+(1-b)(1-q)\\ \frac{b(1-p)}{b(1-p)+(1-b)q}\eqcolon x_{0}\ w.p.\ b(1-p)+(1-b)q.\end{cases}$
Plugging this into the definition of soft accuracy yields
$\operatorname*{\mathrm{SO}}(R(\tilde{s}))$ $=(bp+(1-b)(1-q))((x_{1})^{2}+(1-x_{1})^{2})$
$+(b(1-p)+(1-b)q)((x_{0})^{2}+(1-x_{0})^{2})$
We want to show that $(1-\operatorname*{\mathrm{SO}})-\epsilon(1-b)\geq 0$. We begin by inserting the values, obtaining
$(1-\operatorname*{\mathrm{SO}})-\epsilon(1-b)$ $=\frac{\left(\epsilon\left(b-1\right)+1\right)\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(b\left(p-1\right)+q\left(b-1\right)\right)}{\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(b\left(p-1\right)+q\left(b-1\right)\right)}$
$+\frac{\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(b^{2}\left(p-1\right)^{2}+q^{2}\left(b-1\right)^{2}\right)}{\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(b\left(p-1\right)+q\left(b-1\right)\right)}$
$-\frac{\left(b\left(p-1\right)+q\left(b-1\right)\right)\left(b^{2}p^{2}+\left(b-1\right)^{2}\left(q-1\right)^{2}\right)}{\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(b\left(p-1\right)+q\left(b-1\right)\right)}.$
It is easy to see that the denominator is negative, so it is sufficient to show that the numerator $\operatorname*{\mathrm{NU}}$ is negative. We obtain
$\operatorname*{\mathrm{NU}}$ $=\left(b-1\right)\left(\epsilon b^{2}p^{2}+2\epsilon b^{2}pq-2\epsilon b^{2}p+\epsilon b^{2}q^{2}-2\epsilon b^{2}q+\epsilon b^{2}-2\epsilon bpq+\epsilon bp-2\epsilon bq^{2}\right)$
$+\left(b-1\right)\left(3\epsilon bq-\epsilon b+\epsilon q^{2}-\epsilon q-2b^{2}p^{2}+2b^{2}p+2b^{2}q^{2}-2b^{2}q-2bq^{2}+2bq\right).$
As $b-1$ is negative, it is sufficient to show that $\frac{\operatorname*{\mathrm{NU}}}{b-1}$ is positive. We take the derivative with respect to $\epsilon$:
$\frac{d}{d\epsilon}(\frac{\operatorname*{\mathrm{NU}}}{b-1})=\left(bp+bq-b-q\right)\left(bp+bq-b-q+1\right).$
But the first term equals $b(p-1)+(b-1)q=-\operatorname*{\mathbb{P}}(\tilde{s}=0)$ and is thus between minus one and zero. Correspondingly, the second term is positive, such that the derivative is negative overall. This means that $(\frac{\operatorname*{\mathrm{NU}}}{b-1})$ is minimized at $\epsilon=1$. We insert that value and take two derivatives with respect to $q$:
$\frac{d}{dq}\frac{d}{dq}(\frac{\operatorname*{\mathrm{NU}}}{b-1})=6b^{2}-8b+2.$
This has two zeros at $b=\frac{1}{3}$ and $b=1$ and is negative between them. Correspondingly, under our assumption of $b\geq 0.5$, $\frac{\operatorname*{\mathrm{NU}}}{b-1}$ is concave in $q$ and thus minimized at either of the two extrema $q=0$ or $q=1$. At $q=0,$ we have
$\frac{\operatorname*{\mathrm{NU}}}{b-1}=-b\left(p-1\right)\left(bp+b-1\right).$
This is positive whenever
$bp+b-1\geq 0.$
But at $q=0$, we have $\operatorname*{\mathrm{AG}}=bp$ such that the condition becomes
$\operatorname*{\mathrm{AG}}+b-1\geq 0,$
which is true because we assumed both $\operatorname*{\mathrm{AG}}$ and $b$ to be at least $0.5$.
Meanwhile, At $q=1,$ we have
$\frac{\operatorname*{\mathrm{NU}}}{b-1}=\displaystyle-bp\left(bp-2b+1\right).$
Here it is sufficient to show that
$bp-2b+1\leq 0.$
This time, $\operatorname*{\mathrm{AG}}=bp-b+1$, such that the condition becomes
$\operatorname*{\mathrm{AG}}-b\leq 0,$
which follows from $\epsilon(1-b)\leq 1-\operatorname*{\mathrm{AG}}$, as $\epsilon$ was fixed to one.
C.4 Theorem 12